# Pricing energy spread options with variance gamma-driven Ornstein-Uhlenbeck dynamics

### v2

@@ 7 Concluding remarks
The framework for pricing energy derivatives presented here has wider applicability beyond spread options. Following the explanations in Eberlein, Glau and Papapantoleon EbGlPa2010 and KLW11 ; AlSc2018 , the FFT method can similarly be applied to other payoff functions with a known analytical Fourier transform. Moreover, clearly any European-style payoff in our model can be priced with the Monte Carlo method. Other extensions, such as generalizing the speed of mean reversion $\lambda$ to a matrix, should also be possible, but are beyond the scope of this paper.
Appendix
Appendix A Choice of model parameters
The parameters in Section 5.1 are chosen to approximately resemble two examples of real data on energy prices appearing in the literature. The first marginal component is based on the real data example in (benthbook, , Section 5.1), where $S$ satisfies (4.1) with $\log S(t)=\Lambda(t)+X(t)$, where $X$ is a LDOUP ${\rm d}X(t)-\lambda X(t){\rm d}t+{\rm d}\widetilde{Z}(t)$. The speed of mean reversion is $\widetilde{\lambda}=0.073$ under their time scale where $t=250$ represents 1 year, and changing to our time scale where $t=1$ does implies $\lambda=\widetilde{\lambda}/\Delta$, where $\Delta=\frac{1}{250}$, with no change to the distribution of $\widetilde{Z}$. This follows from the fact that if ${\bf X}\sim{OU}(\lambda,{\bf Z})$, then $({\bf X}(ct))_{t\geq 0}\sim{OU}(c\lambda,{\bf Z})$ for $c>0$. Thus, we obtain $\lambda$ in (5.5).
The BDLP is a normal inverse Gaussian process $\widetilde{Z}\sim\operatorname{NIG}^{1}(\alpha,\beta,\gamma,\eta)$ with the usual parameterization (see (sco, , Section 5.3.8), $\eta$ here is $\mu$ there). Specifically, their model has $(\alpha,\beta,\gamma)=(4.869,-0.125,0.0874)$, and we set $\eta=0.00284$ to make $\mathbb{E}[\widetilde{Z}(1)]=0$. The VG distribution matching the 2nd, 3rd, 4th central moments is $\widetilde{Z}\sim VG^{1}(0.42485,\allowbreak-0.00224,0.01796)$. Finally, noting that $\widetilde{Z}=Z(\lambda t)$, we obtain the parameters $\alpha_{1},\mu_{1},\Sigma_{11}$ in (5.5). Similarly, preliminary seasonality parameters $(b_{01},\dots,b_{31})$ can be obtained to exactly match theirs, accounting for the change in time scale.
The second marginal component is based on the real data example in bs14 . This model is $S(t)=\Lambda(t)+Y(t)+X(t)$, where $Y$ is a NIG process and $X$ is a OU-NIG process like in the first marginal component, where both NIG processes are independent. In particular, it is not of the form (4.1) and is an arithmetic model.
We simulate this model but deviate from bs14 in a few ways. We change from their time scale where $t=365$ represents 1 year to ours where $t=1$ does, and take the sampling interval to be $\Delta=\frac{1}{250}$ as above. We set their parameter $\xi_{0}=40$, rather than 738.77, since this is approximately consistent with their plotted seasonality function, and set $\xi_{4}=\xi_{5}=0$ to remove the weekend indicators. The simulation has 1001 observations, and is simulated until, like the real data, a completely positive sample path is obtained. Then we fit the marginal components of our model in Section 4.1 to this to obtain the parameters $\alpha_{2},\mu_{2},\Sigma_{22}$ in (5.5) and obtain preliminary estimates of the seasonal parameters $(b_{20},\dots,b_{23})$. Since the bivariate model for $\bf X$} has only one $\lambda$, we use the value obtained from the first marginal component.
Energy prices tend to have very high correlation. Accordingly, the joint parameters $a,\Sigma_{12}$ are chosen so that the Brownian motion correlation is $\rho={\Sigma_{12}}/\sqrt{\Sigma_{11}\Sigma_{22}}=0.95$, and $a=0.9(\frac{1}{\alpha_{1}}\wedge\frac{1}{\alpha_{2}})$. Both are close to the maximum permitted by the parameter constraint $\rho\in(-1,1)$ and $a\in(0,(\frac{1}{\alpha_{1}}\wedge\frac{1}{\alpha_{2}}))$, so that $\operatorname{Corr}(Z_{1}(1),Z_{2}(1))=0.50303$ close to the maximum subject to the marginal components specified above. To obtain a correlation approximately 1 for a WVAG process $\bf Z$} would require both $Z_{1}$ and $Z_{2}$ to be approximately equal in law.
Finally, note that the cyclical component of the seasonality function $\Lambda_{k}(t)$ can be converted into the form
$b_{2k}\cos\mathopen{}\left(\frac{2\pi}{p}t\right)\mathclose{}+b_{3k}\sin\mathopen{}\left(\frac{2\pi}{p}t\right)\mathclose{}=A\cos\mathopen{}\left(\frac{2\pi(t-\phi)}{p}\right)\mathclose{},$
where $A$ is the amplitude and $\phi\in[0,1]$ is the phase. The phase is arbitrary and it is reasonable, but not necessary, to expect both energy prices have the same phase, so we convert the preliminary seasonal parameters to have $\phi=0.22$ while keeping $A$ unchanged. The intercept $(b_{01},b_{02})$ affects the price level and may also be considered arbitrary. We set it such that $\boldsymbol{\Lambda}(T_{e})=\log{\bf S}(T_{e})=\log(100,96)$, which is the current energy price. This choice corresponds to assuming ${\bf S}(T_{e})$ is such that ${\bf X}(T_{e})$ is at its stationary mean, with different choices of $(b_{01},b_{02})$ corresponding to assuming it is above or below that mean. This leads to seasonal parameters in (5.1)–(5.2).
The MPR being set to ${\bf h}=(-0.1,-0.03)$ is arbitrary but appears to be within the range of plausible values. See, for example, (BenthPir18, , Section 4.1) in the context of wind power futures.
@@ Acknowledgments
Kevin Lu thanks Mesias Alfeus for discussions relating to this work.

### v3

@@ 8 Concluding remarks
The framework for pricing energy derivatives presented here has wider applicability beyond spread options. Following the explanations in Eberlein, Glau and Papapantoleon EbGlPa2010 and KLW11 ; AlSc2018 , the FFT method can similarly be applied to other payoff functions with a known analytical Fourier transform. Moreover, any European-style payoff in our model can readily be priced using the Monte Carlo method. Other extensions, such as generalizing the speed of mean reversion $\lambda$ to a matrix, using a generalized hyperbolic (GH) process, or other processes as the BDLP, would likely lead to a better fit in Figure 8 because the GH process has 1 additional shape parameter and includes the VG process as a subclass. Another direction is to consider an arithmetic model. However, these are beyond the scope of this paper.
@@ Acknowledgments
Kevin Lu thanks Mesias Alfeus for discussions relating to this work. We thank the referees for their comments.